OpenStudy (anonymous):

combine and simplify fraction (1/1-sinx)+(1/1+sinx)

OpenStudy (abb0t):

Find the common denominator between the two denominators and combine. For example: \[\frac{ 1 }{ A } + \frac{ 2 }{ B } = \frac{ B+2A }{ AB }\]

OpenStudy (anonymous):

thanks!